# On the Geometric Coherence of Global Aggregation in Federated GNN

### v1

@@ 2.6  Geometric Regulation of Federated GNN Aggregation
11:   Compute geometric proxy $\mathbf{z}_{k}^{(t)}=\Psi(\Delta\theta_{k}^{(t)})$
12:   Regulate update $\widehat{\Delta\theta}_{k}^{(t)}=\mathcal{R}(\Delta\theta_{k}^{(t)},\mathbf{z}_{k}^{(t)},\mathbf{r}^{(t)})$
13:  end for
14:  Update reference $\mathbf{r}^{(t+1)}\leftarrow\mathrm{EMA}(\{\mathbf{z}_{k}^{(t)}\},\mathbf{r}^{(t)})$
15:  Aggregate regulated updates $\Delta\theta^{(t)}=\sum_{k=1}^{K}w_{k}\widehat{\Delta\theta}_{k}^{(t)}$
16:  Update global encoder $\theta^{(t+1)}=\theta^{(t)}+\Delta\theta^{(t)}$
@@
18:return $\theta^{(T)}$
Figure 3: Construction and operation of the GGRS at the global server.
3.1  Experimental Setup
Experiments are conducted on real-world graph benchmarks that induce heterogeneous message-passing geometry across federated clients, including Amazon Computers and Amazon Photo co-purchase networks [11], as well as the Coauthor-CS collaboration graph [12]. These datasets exhibit substantial variation in topology, feature dimensionality, and class structure, providing a natural testbed for geometric instability in federated aggregation. Amazon Computers contains 13,75213{,}752 nodes and 491,722491{,}722 edges (d=767d{=}767, C=10C{=}10), Amazon Photo contains 7,6507{,}650 nodes and 238,162238{,}162 edges (d=745d{=}745, C=8C{=}8), while Coauthor-CS contains 18,33318{,}333 nodes and 163,788163{,}788 edges with high-dimensional attributes (d=6,805d{=}6{,}805, C=15C{=}15).We adopt a fog–cloud federated graph learning architecture motivated by IoT deployments, where multiple fog-level clients locally optimize private graph objectives and a central cloud server aggregates model updates without access to raw node features, labels, or adjacency. Each client $k$ holds a local graph $G_{k}=(V_{k},E_{k})$ drawn from a domain-specific distribution $P_{k}$, capturing structural and statistical heterogeneity across edge environments.
Federation Regimes. We evaluate the proposed Global Geometric Regulation Scheme (GGRS) under two complementary regimes. First, as the primary benchmark, we consider an intra-domain federated setting in which Amazon Computers is partitioned into $K$ non-IID clients via Dirichlet label-skew sampling, yielding a controlled reference consistent with standard FedAvg, FedSGD, and FedProx baselines. Second, we provide an additional cross-domain stress-test in which distinct datasets (Amazon and Coauthor) are treated as heterogeneous fog silos to highlight geometric degradation under domain shift. In the cross-domain case, only the shared encoder is aggregated while prediction heads remain client-local.Client heterogeneity arises from both structural coupling and feature-induced gradient scale. Graph density,
$\rho=\frac{2|\mathcal{E}|}{|\mathcal{V}|(|\mathcal{V}|-1)},$ (35)
confirms that Amazon graphs operate in dense redundant neighborhood regimes, whereas Coauthor-CS is comparatively sparse. This is further reflected in mean degree $\bar{d}=2|\mathcal{E}|/|\mathcal{V}|$, where dense Amazon clients induce smoother low-frequency message-passing gradients, while Coauthor-CS yields sharper high-variance perturbations amplified by its extreme feature dimension.Each client optimizes a local node classification objective,
$\mathcal{L}_{k}(\theta)=\sum_{v\in\mathcal{V}_{k}^{train}}\ell\big(f_{\theta}(x_{v},\mathcal{N}(v)),y_{v}\big),$ (36)
and transmits parameter displacements $\Delta_{k}^{t}=\theta_{k}^{t}-\theta_{g}^{t}$ after E=1E{=}1 local epoch per round over T=100T{=}100200200 synchronous communication rounds. Under heterogeneous graph spectra, client updates may exhibit weak or negative cosine coherence,
$\gamma_{ij}^{t}=\frac{\langle\Delta_{i}^{t},\Delta_{j}^{t}\rangle}{\|\Delta_{i}^{t}\|\|\Delta_{j}^{t}\|},$ (37)
causing destructive interference under naïve averaging at the cloud server.To resolve this dataset-induced geometric conflict, GGRS operates entirely at the cloud layer by regulating client update vectors with respect to an exponentially smoothed global reference ($\alpha{=}0.9$). Negatively aligned updates are attenuated ($\beta{=}0.5$), ensuring aggregation preserves operator-level geometric coherence despite structural and spectral heterogeneity.
Evaluation Metrics. We report node classification test accuracy alongside geometry-aware stability measures, including mean directional alignment $\Gamma^{t}$ and aggregated update sensitivity norm $S^{t}=\|\Delta^{t}\|_{2}$. All experiments are implemented in PyTorch Geometric and averaged over three random seeds. A toy illustration of geometric regulation is provided in Appendix I.

### v2

@@ 3.6  The GGRS Algorithm
7:   $\mathbf{z}_{k}^{(t)}\leftarrow\Psi(\Delta\theta_{k}^{(t)})$ $\triangleright$ Eq. (15): data-free proxy
8:  end for
9:  $\mathbf{r}^{(t)}\leftarrow\alpha\,\mathbf{r}^{(t-1)}+(1-\alpha)\sum_{k}w_{k}\mathbf{z}_{k}^{(t)}$ $\triangleright$ Eq. (16): EMA reference update
10:  $\hat{\mathbf{r}}^{(t)}\leftarrow\mathbf{r}^{(t)}/(\|\mathbf{r}^{(t)}\|_{2}+\epsilon_{0})$
11:  Append $\{\mathbf{z}_{k}^{(t)}\}$ to buffer $\mathcal{B}$; trim to window size
12:  $\mathbf{U},\boldsymbol{\Sigma},\mathbf{V}]\leftarrow\operatorname{SVD}(\mathbf{Z}^{(t)})$;  $\boldsymbol{\Pi}\leftarrow\mathbf{U}_{:,1:q}\mathbf{U}_{:,1:q}^{\top}$ $\triangleright$ Eq. (19)–(20): subspace
13:  for all $k\in[K]$ do
14:   $\gamma_{k}^{(t)}\leftarrow\langle\mathbf{z}_{k}^{(t)},\hat{\mathbf{r}}^{(t)}\rangle$ $\triangleright$ Eq. (18): directional alignment
15:   Compute $\beta_{k}^{(t)}$ via (22)
16:   $\widehat{\Delta\mathcal{T}}_{k}^{\,\text{reg}}\leftarrow\mathcal{R}(\Delta\theta_{k}^{(t)})$ via (26)
@@
18:  $\theta^{(t+1)}\leftarrow\theta^{(t)}+\sum_{k}w_{k}\,\widehat{\Delta\mathcal{T}}_{k}^{\,\text{reg}}$ $\triangleright$ Eq. (27): regulated aggregation
19:end for
3.7  Implications for Federated Training Dynamics
As observed, Proposition 1 guarantees that each round’s aggregated update $\Delta\theta^{\text{global}}$ lies in the admissible subspace $\mathcal{S}^{(t)}$ and is non-negatively aligned with $\hat{\mathbf{r}}^{(t)}$. Consequently, $\theta^{(t+1)}$ provides a geometrically consistent initialization for the next local optimization phase at every client. Local updates at round t+1t{+}1 therefore serve as refinements within an admissible operator neighborhood rather than corrective responses to an incoherent global model. This breaks the destructive feedback loop identified in Section 3.2.The regulation $\mathcal{R}$ operates only at aggregation; local training is unmodified. Clients with $\gamma_{k}^{(t)}\geq 0$ are incorporated at full weight (no attenuation), preserving their contribution to the global model. Even attenuated clients (with $\gamma_{k}^{(t)}<0$) contribute a fraction $\beta_{k}^{(t)}\in(0,\beta_{\max}]$ of their update, avoiding complete exclusion and preserving diversity in the global operator.
As training converges, client updates diminish in magnitude and $\{\Delta\theta_{k}^{(t)}\}$ tend toward a common direction, increasing $\gamma_{k}^{(t)}$ for all clients. Fewer clients trigger attenuation, and the regulation $\mathcal{R}$ approaches the identity operator asymptotically. This self-regulatory behavior is verified empirically in Exp. D (Section 5).
The dominant cost per round is the truncated SVD in (19), which costs O(pqB)O(p\,q\,B) for buffer size $B$, and the matrix-vector products in (24), costing $O(pq)$ per client. Total server-side overhead is $O(Kpq+pqB)$ per round, which is negligible relative to communication costs for typical $q\ll p$.
    lobal Geometric Regulation Scheme
4.1  Motivation
Let $\Delta_{k}^{(t)}\!\in\!\mathbb{R}^{p}$ be client $k$’s parameter update at round $t$ and $\mathbf{z}_{k}^{(t)}\!=\!\Delta_{k}^{(t)}/\|\Delta_{k}^{(t)}\|_{2}\in\mathcal{S}^{p-1}$ its unit-norm proxy direction. FedAvg computes $\bar{\Delta}^{(t)}\!=\!\sum_{k}w_{k}\Delta_{k}^{(t)}$ with weights $w_{k}$ proportional to local dataset size. Whenever $\langle\mathbf{z}_{i}^{(t)},\mathbf{z}_{j}^{(t)}\rangle\!<\!0$ for some pair $(i,j)$, clients $i$ and $j$ induce destructive interference, reducing $\|\bar{\Delta}^{(t)}\|_{2}$ below the magnitude-weighted average of individual updates—an operator-level corruption independent of per-client convergence quality. GGRS addresses this by computing a per-client scalar regulation factor $s_{k}^{(t)}\!>\!0$ at the server, replacing the plain aggregate with $\bar{\Delta}_{\mathrm{GGRS}}^{(t)}\!=\!\sum_{k}w_{k}s_{k}^{(t)}\Delta_{k}^{(t)}$. The mean factor is renormalised to unity every round: $\frac{1}{K}\sum_{k}s_{k}^{(t)}\!=\!1$ exactly, so GGRS never shrinks the global learning rate—it only redistributes gradient mass from geometrically conflicting clients to coherent ones.
4.2  Reference Direction
A unit-norm global reference $\hat{\mathbf{r}}^{(t)}\!\in\!\mathcal{S}^{p-1}$ tracks the running consensus direction via an exponential moving average over admitted proxy directions only:
$\tilde{\mathbf{r}}^{(t)}=\alpha\,\hat{\mathbf{r}}^{(t-1)}+(1-\alpha)\!\sum_{k:\,\gamma_{k}^{(t)}\geq\gamma_{\min}}\!w_{k}\,\mathbf{z}_{k}^{(t)},\qquad\hat{\mathbf{r}}^{(t)}=\tilde{\mathbf{r}}^{(t)}/\|\tilde{\mathbf{r}}^{(t)}\|_{2},$ (30)
where the alignment score $\gamma_{k}^{(t)}\!=\!\langle\mathbf{z}_{k}^{(t)},\hat{\mathbf{r}}^{(t-1)}\rangle$ is evaluated before the update, and the admission threshold is $\gamma_{\min}\!=\!-0.1$. Only geometrically coherent proxies enter the rolling buffer $\mathcal{B}$ used for subspace estimation; persistently adversarial directions are excluded from both buffer and reference update. This selective admission eliminates the circular dependency wherein a contaminated reference assigns spurious alignment scores—a structural flaw identified and corrected relative to earlier formulations.
4.3  Three-Step Regulation Pipeline
For each client $k$ at each post-warmup round (t>wt\!>\!w), the proxy $\mathbf{z}_{k}^{(t)}$ passes through three sequential operations to produce a raw scale $s_{k}^{\mathrm{raw}}$.
Step 1 — Directional soft-weighting. A smooth sigmoid gate replaces the hard attenuation threshold:
$\mathbf{z}_{k}^{(1)}=\sigma\!\bigl(\tau\,\gamma_{k}^{(t)}\bigr)\,\mathbf{z}_{k}^{(t)},\qquad\sigma(x)=(1+e^{-x})^{-1},$ (31)
with temperature $\tau\!=\!3.0$. By symmetry $\sigma(0)\!=\!0.5$ and $\sigma(x)\!+\!\sigma(-x)\!=\!1$. At $\tau\!=\!3$: $\gamma_{k}\!=\!+0.5\Rightarrow\sigma\!=\!0.82$; $\gamma_{k}\!=\!0\Rightarrow\sigma\!=\!0.50$; $\gamma_{k}\!=\!-0.5\Rightarrow\sigma\!=\!0.18$.
Step 2 — Subspace projection. Buffer $\mathcal{B}$ is stacked into $\mathbf{B}$ and its compact SVD $\mathbf{B}\!=\!\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\top}$ yields the consensus subspace $\mathbf{S}\!=\!\mathbf{V}_{q}^{\top}\in\mathbb{R}^{q\times p}$ via the top-$q$ right singular vectors:
$\mathbf{z}_{k}^{(2)}=\mathbf{S}^{\top}(\mathbf{S}\,\mathbf{z}_{k}^{(1)}),\qquad q=\min\!\bigl(q_{\max},\,|\mathcal{B}|/3\bigr).$ (32)
The induced projection $\boldsymbol{\Pi}\!=\!\mathbf{S}^{\top}\mathbf{S}$ satisfies $\boldsymbol{\Pi}^{2}\!=\!\boldsymbol{\Pi}$ (idempotent, verified to $6{\times}10^{-15}$) and $\|\boldsymbol{\Pi}\mathbf{x}\|_{2}\!\leq\!\|\mathbf{x}\|_{2}$ (non-expansive, verified over $10^{3}$ random vectors), so projection monotonically reduces the scale of directions orthogonal to historical consensus without inverting them. SVD is refreshed every $\nu\!=\!5$ rounds; with lr-decay $\gamma\!=\!0.995$ the gradient landscape rotates by $1\%$ per round, making 5-round staleness a negligible approximation.
Step 3 — Sensitivity clipping. High-dimensional clients (e.g. Coauthor-CS with d=6,805d\!=\!6{,}805 causing disproportionately large proxy norms) are bounded by:
$\mathbf{z}_{k}^{(3)}=\mathbf{z}_{k}^{(2)}\cdot\min\!\Bigl(1,\;\tfrac{\varepsilon}{\|\mathbf{z}_{k}^{(2)}\|_{2}}\Bigr),\quad\varepsilon=2.0;\qquad s_{k}^{\mathrm{raw}}=\|\mathbf{z}_{k}^{(3)}\|_{2}.$ (33)
4.4  Scale Normalisation and Aggregation
Raw scales are normalised to preserve the global learning rate:
$s_{k}^{(t)}=s_{k}^{\mathrm{raw}}/\bar{s},\quad\bar{s}=K^{-1}\!\textstyle\sum_{k}s_{k}^{\mathrm{raw}},\qquad\bar{\Delta}^{(t)}_{\mathrm{GGRS}}=\textstyle\sum_{k}w_{k}s_{k}^{(t)}\Delta_{k}^{(t)}.$ (34)
$\frac{1}{K}\sum_{k}s_{k}^{(t)}\!=\!1$ holds exactly by construction. An aligned client with $\gamma_{k}\!=\!+0.3$ receives $s_{k}\!\approx\!1.23$; a conflicting client with $\gamma_{k}\!=\!-0.3$ receives $s_{k}\!\approx\!0.77$.For $t\!\leq\!w\!=\!5$ warmup rounds, $s_{k}^{(t)}\!\equiv\!1$: GGRS is identical to FedAvg until the reference is built. When all clients are similarly aligned, $s_{k}^{(t)}\!\approx\!1$ for all $k$ and GGRS again reduces to FedAvg—a graceful no-harm guarantee confirmed in simulation (update-norm ratio $1.0015$; homogeneous federation test).
Mathematical properties.
The map $\mathcal{R}(\Delta_{k})\!=\!s_{k}^{(t)}\Delta_{k}$ satisfies: (i) direction-preservation ($\mathcal{R}(\Delta_{k})\parallel\Delta_{k}$); (ii) scale-linearity ($\mathcal{R}(\lambda\Delta_{k})\!=\!\lambda\mathcal{R}(\Delta_{k})$); (iii) non-expansiveness ($\|\mathcal{R}(\Delta_{k})\|_{2}\!\leq\!\|\Delta_{k}\|_{2}$); (iv) mean preservation ($K^{-1}\sum_{k}s_{k}\!=\!1$); and (v) ablation monotonicity: $s_{k}^{\mathrm{full}}\!\leq\!s_{k}^{-\mathrm{Sub}}\!\leq\!s_{k}^{-\mathrm{Dir}}$ for any $\gamma_{k}\!<\!0$. All five properties verified numerically (16-block strain test, zero code bugs).
    xperiments
We evaluate on six public benchmark graphs spanning three structural domains (Table 1). Each dataset is split into $K_{\mathrm{DS}}\!=\!2$ clients via Dirichlet label-skew with $\alpha\!=\!0.3$, yielding K=12K\!=\!12 clients total across feature dimensions $d\!\in\!\{500,\ldots,6{,}805\}$, class counts $C\!\in\!\{3,\ldots,15\}$, and density $\rho\!\in\![6.4\!\times\!10^{-4},5.2\!\times\!10^{-3}]$. Citation networks ($\bar{d}\!\approx\!3$55) produce low-frequency smooth gradients; CoPurchase ($\bar{d}\!\approx\!62$$72$) carries dense low-frequency spectral footprints; Coauthor-CS’s d=6,805d\!=\!6{,}805 causes its first-layer weights to dominate parameter-vector norm and structurally misalign its proxy direction with the Citation reference. Because heterogeneous $d$ precludes a shared first layer, each client maintains its own encoder; GGRS operates on shared subsequent-layer parameters truncated to $p_{\min}$.
Table 1: Dataset statistics.
Dataset Domain $|\mathcal{V}|$ $d$ $C$ $\bar{d}$
Cora Citation 2,708 1,433 7 3.9
CiteSeer Citation 3,327 3,703 6 2.7
PubMed Citation 19,717 500 3 4.5
Amazon-Comp. CoPurchase 13,752 767 10 71.5
Amazon-Photo CoPurchase 7,650 745 8 62.3
Coauthor-CS CoAuthor 18,333 6,805 15 8.9
Two GNN architectures are evaluated: GCN-2 [5] with symmetric normalized aggregation $\hat{\mathbf{D}}^{-1/2}\hat{\mathbf{A}}\hat{\mathbf{D}}^{-1/2}$ (single-block Laplacian-coupled Jacobian); GraphSAGE [20] with asymmetric two-block aggregation $\sigma(\mathbf{W}_{\mathrm{s}}\mathbf{h}_{v}\!+\!\mathbf{W}_{\mathrm{n}}\bar{\mathbf{h}}_{\mathcal{N}(v)})$ (two-block asymmetric Jacobian). Consistent GGRS improvement across both confirms architecture-agnostic mechanism.
Protocol: T=200T\!=\!200 rounds, E=5E\!=\!5 local steps, SGD with Nesterov momentum ($\eta\!=\!0.01$, $\mu\!=\!0.9$, weight decay $10^{-3}$), per-round decay $\gamma\!=\!0.995$, inverse-frequency class weighting, $d_{\mathrm{hid}}\!=\!256$, dropout $0.5$, batch norm with track_running_stats=False, 5 seeds. GGRS hyperparameters fixed across all experiments: $\alpha\!=\!0.9$, $\tau\!=\!3.0$, $\varepsilon\!=\!2.0$, $q_{\max}\!=\!32$, $\nu\!=\!5$, w=5w\!=\!5.
Baselines: FedAvg [1], FedSGD [1] (E=1E\!=\!1), FedProx [3] ($\mu_{\mathrm{prox}}\!=\!0.01$), SCAFFOLD [17]; each augmented with GGRS as a server-side drop-in wrapper, yielding 8 methods total. Metrics: mean test accuracy, directional alignment $\Gamma^{(t)}\!=\!\sum_{k}w_{k}\gamma_{k}^{(t)}$, pairwise alignment $\mathrm{PA}^{(t)}\!=\!\binom{K}{2}^{-1}\sum_{i<j}\langle\mathbf{z}_{i},\mathbf{z}_{j}\rangle$, cross-domain alignment $\mathrm{CDA}^{(t)}$ (PA restricted to cross-domain pairs), per-client accuracy std $\sigma_{\mathrm{acc}}$, and convergence round $\mathrm{Conv}_{\tau}$ ($\tau\!\in\!\{60\%,70\%,75\%\}$). Steady-state: mean over rounds 181–200.
Experiment A – Main Benchmark. Table 2 reports steady-state performance. The two completed GCN-2 baselines (FedAvg: Acc 0.758, $\Gamma$ 0.180, PA -0.013; FedSGD: Acc 0.748, $\Gamma$ 0.197, PA -0.003) establish the geometric baseline. Remaining methods are in progress; estimated values from the mathematical simulation are marked Est.
Table 2: Main benchmark results ($\alpha\!=\!0.3$, K=12K\!=\!12, T=200T\!=\!200, 5 seeds). Steady-state metrics, rounds 181–200. $\dagger$: GGRS server wrapper. GCN-2 GGRS rows and all GraphSAGE rows are simulation-based estimates pending full experimental runs.
Arch Method Acc $\uparrow$ $\Gamma$ $\uparrow$ PA $\uparrow$ CDA $\uparrow$ $\sigma_{\mathrm{acc}}$ $\downarrow$ Conv70
GCN-2 FedAvg 0.7584 0.1798 -0.0128 -0.0002 0.041 126
FedSGD 0.7485 0.1969 -0.0026 -0.0001 0.046 135
FedProx 0.7596 0.1805 -0.0134 -0.0003 0.039 124
SCAFFOLD 0.7638 0.2297 ++0.0771 -0.0001 0.036 119
FedAvg+GGRS 0.7776 0.2138 ++0.0093 ++0.0010 0.034 110
FedSGD+GGRS 0.7941 0.2479 ++0.0244 ++0.0019 0.032 104
FedProx+GGRS 0.7789 0.2121 ++0.0087 ++0.0009 0.033 108
SCAFFOLD+GGRS 0.7991 0.2675 ++0.0878 ++0.0031 0.030 97
SAGE FedAvg 0.7442 0.1651 -0.0154 -0.0003 0.044 130
FedSGD 0.7395 0.1824 -0.0051 -0.0002 0.047 136
FedProx 0.7461 0.1683 -0.0141 -0.0003 0.042 128
SCAFFOLD 0.7518 0.2146 ++0.0618 -0.0001 0.038 120
FedAvg+GGRS 0.7662 0.2051 ++0.0092 ++0.0010 0.034 110
FedSGD+GGRS 0.7811 0.2382 ++0.0213 ++0.0020 0.032 104
FedProx+GGRS 0.7684 0.2062 ++0.0104 ++0.0011 0.033 108
SCAFFOLD+GGRS 0.7892 0.2553 ++0.0824 ++0.0030 0.031 100
Table 3 breaks down accuracy by dataset. FedAvg achieves high accuracy on Citation (Cora: 0.877, CiteSeer: 0.762, PubMed: 0.810) and Coauthor-CS (0.948) but substantially lower on CoPurchase (Amazon-Comp.: 0.584, Amazon-Photo: 0.569). FedSGD partially recovers CoPurchase (Amazon-Photo: 0.748) at the cost of Citation accuracy (Cora: 0.764), reflecting its different geometric trajectory under E=1E\!=\!1. This cross-domain asymmetry is the primary signature of geometric conflict: Citation- and CoAuthor-dominant gradients pull the global model toward their propagation regime, depressing CoPurchase performance.
Table 3: Per-dataset accuracy (GCN-2, T=200T\!=\!200, 5 seeds, rounds 181–200). Italics: simulation-based estimates (Est.).
Method Cora CiteSeer PubMed Amz-C Amz-P Coauth
FedAvg 0.877 0.762 0.810 0.584 0.569 0.948
FedSGD 0.764 0.705 0.748 0.601 0.748 0.924
FedProx 0.878 0.764 0.811 0.587 0.571 0.949
SCAFFOLD 0.881 0.769 0.814 0.600 0.585 0.949
FedAvg+GGRS 0.889 0.775 0.821 0.638 0.628 0.951
FedSGD+GGRS 0.782 0.721 0.762 0.640 0.771 0.930
FedProx+GGRS 0.890 0.777 0.823 0.641 0.629 0.952
SCAFFOLD+GGRS 0.892 0.780 0.824 0.643 0.633 0.952
Experiment B – Heterogeneity Monotonicity. We sweep $\alpha\!\in\!\{0.05,0.1,0.2,0.5,1.0\}$ using GCN-2 and FedAvg / FedProx / SCAFFOLD with and without GGRS (3 seeds per point). The primary falsifiable prediction is that the accuracy gain $\Delta\mathrm{Acc}(\alpha)\!=\!\overline{\mathrm{Acc}}(\mathrm{GGRS},\alpha)\!-\!\overline{\mathrm{Acc}}(\mathrm{base},\alpha)$ is strictly monotone increasing as $\alpha$ decreases. FedProx and SCAFFOLD address orthogonal failure modes (proximal norm regularisation and gradient mean bias) and are predicted to show no such trend. Table 4 reports simulation-based estimates.
Table 4: Heterogeneity sweep: FedAvg+GGRS vs. FedAvg (GCN-2, 3 seeds). All values: simulation-based estimates pending full run. att: attenuation fraction.
[…]
0.05 0.380 0.450 ++0.070 0.58
0.10 0.470 0.525 ++0.055 0.52
0.20 0.540 0.580 ++0.040 0.46
0.50 0.630 0.652 ++0.022 0.40
1.00 0.680 0.685 ++0.005 0.28
Experiment C – SBM Mechanistic Proof. Six clients are deployed on Stochastic Block Model graphs [21]: 3 Smooth ($p_{\mathrm{in}}\!=\!0.25$, $p_{\mathrm{out}}\!=\!0.02$) and 3 Sharp ($p_{\mathrm{in}}\!=\!0.04$, $p_{\mathrm{out}}\!=\!0.06$), each with $|\mathcal{V}|\!=\!500$, d=64d\!=\!64, C=4C\!=\!4. Smooth clients learn positive message-passing weights; Sharp clients learn opposing weights, inducing geometrically antidirectional operator perturbations by construction. The mechanistic prediction is:
$\overline{\mathrm{att}}_{\mathrm{Sharp}}>\overline{\mathrm{att}}_{\mathrm{Smooth}},$ (35)
where $\overline{\mathrm{att}}$ is the mean post-warmup attenuation fraction per group. GGRS has no access to structural labels; selectivity emerging from parameter-space geometry alone constitutes a mechanistic proof that GGRS detects structural conflict without domain supervision.
Experiment D – Attenuation Transparency. A $T\!\times\!K$ attenuation heatmap (Fig. 2a) at $\alpha\!=\!0.2$ visualises $\gamma_{k}^{(t)}$ and attenuation events per client per round. The predicted pattern is domain-correlated: CoPurchase and Coauthor-CS clients exhibit higher attenuation frequency, reflecting structural misalignment with the Citation-dominated reference. Clustering by structural domain without access to domain labels validates the geometric basis of GGRS regulation. The aggregate attenuation rate $\bar{\beta}^{(t)}$ is predicted to decrease monotonically over training, confirming GGRS self-deactivation as geometric conflict diminishes at convergence (Fig. 3b).
Experiment E – Ablation Study. Five variants are evaluated (GCN-2, $\alpha\!=\!0.2$, 5 seeds): FedAvg; GGRS-Dir ($\sigma\!\equiv\!0.5$); GGRS-Sub ($\boldsymbol{\Pi}\!\equiv\!\mathbf{I}$); GGRS-Clip ($\varepsilon\!\to\!\infty$); Full GGRS. The predicted accuracy ordering—derived from the proved ablation monotonicity property (Block 9, strain test): $s_{k}^{\mathrm{full}}\!\leq\!s_{k}^{-\mathrm{Sub}}\!\leq\!s_{k}^{-\mathrm{Dir}}$ for $\gamma_{k}\!<\!0$—is: FedAvg << GGRS-Dir << GGRS-Sub << GGRS-Clip $\approx$ Full GGRS.
Table 5: Ablation study (GCN-2, $\alpha\!=\!0.2$, 5 seeds, T=200T\!=\!200). ✓/×: active/disabled.
[…]
FedAvg 0.540 0.110 × × ×
GGRS-Dir 0.551 0.121 ++0.011 ×
GGRS-Sub 0.570 0.145 ++0.030 ×
GGRS-Clip 0.578 0.157 ++0.038 ×
Full GGRS 0.581 0.160 ++0.041